# Large Bayesian Tensor Autoregressions

### v1

@@ 2.2 Introduction to Tensor Autoregressive Models
where $I=I_{1}I_{2}I_{3}$, and $\text{vec}(\mathcal{E}_{t})=\text{vec}(\mathbf{E}_{t,(1)})$, with $\text{vec}(\mathbf{X})$ denoting the operation that stacks the columns of the matrix $\mathbf{X}$ vertically into a single vector. $\boldsymbol{\Sigma}_{1}$, $\boldsymbol{\Sigma}_{2}$, and $\boldsymbol{\Sigma}_{3}$ are $I_{1}\times I_{1}$, $I_{2}\times I_{2}$, and $I_{3}\times I_{3}$ covariance matrices, respectively. The Kronecker structure is adopted to achieve dimensionality reduction. Without this structure, sampling $\boldsymbol{\Sigma}$ would require estimating $\mathcal{O}(I^{2})$ free parameters, making estimation unreliable with limited samples and computationally expensive—or even infeasible—as $I$ increases, especially in high-dimensional tensor time series. In contrast, imposing a Kronecker structure drastically reduces the number of free parameters to $\sum_{i=1}^{6}{I_{i}R_{i}}$ and lowers the computational cost to $\mathcal{O}(\max(I_{1},I_{2},I_{3})^{6})$, thereby making the estimation more reliable and computationally tractable even in high-dimensional settings. To sample $\boldsymbol{\Sigma}_{i}$, we first compute the mode-$i$ matricization of the tensor $\mathcal{E}_{t}$, denoted by $\mathbf{E}_{t,(i)}$. It can be readily shown that $\mathbf{E}_{t,(i)}$ satisfies the following property:
@@ 5.3 Interpretation for Response Factor Time Series
In the BTAR-TK model with shrinkage priors, we adopt a dynamic factor interpretation for both responses and predictors. Specifically, the 6464 $(=R_{1}R_{2}R_{3}=R_{4}R_{5}R_{6})$ lagged response and predictor factor time series, $\{\widetilde{\mathbf{B}}_{m}^{\prime}\mathbf{y}_{t}\}$ and $\{\widetilde{\mathbf{B}}^{\prime}\mathbf{y}_{t-1}\}$, are constructed; in the main text we focus on the response factors. Although the model is estimated on disaggregated country- and commodity-level data, most dimension-reduced response factors closely track more aggregated global indicators; the corresponding indicators are also displayed in the figure. Because each column of $\mathbf{B}_{i}$ is identified only up to sign, the sign of each factor is indeterminate; we orient each factor by choosing the sign that minimizes the RMSE with the corresponding index.
@@ 6 Conclusion
In summary, we introduce a flexible and computationally efficient Bayesian framework for third-order tensor autoregressions based on Tucker decomposition. Our approach’s key innovations are the integration of stochastic volatility and hierarchical shrinkage priors—which automatically shrink negligible tensor coefficients toward zero—into high-dimensional tensor-on-tensor regression. In extensive MCMC experiments, we show that when the true data-generating process is low-rank, the Tucker model, especially the model with multiway stick breaking shrinkage priors, tend to outperform both CP decomposition and standard Bayesian VARs. Even if the true coefficient tensor lacks an assumed low-rank structure, our method remains competitive with Bayesian VARs on small and medium datasets. Finally, in an empirical application to $19\times 19\times 15$ multi-category bilateral trade flows, we identify the principal countries and commodity groups driving global trade dynamics. We also provide a factor interpretation that links the model’s factors to global indicators.

### v2

@@ 2.2 Introduction to Tensor Autoregressive Models
where $I=I_{1}I_{2}I_{3}$, and $\text{vec}(\mathcal{E}_{t})=\text{vec}(\mathbf{E}_{t,(1)})$, with $\text{vec}(\mathbf{X})$ denoting the operation that stacks the columns of the matrix $\mathbf{X}$ vertically into a single vector. $\boldsymbol{\Sigma}_{1}$, $\boldsymbol{\Sigma}_{2}$, and $\boldsymbol{\Sigma}_{3}$ are $I_{1}\times I_{1}$, $I_{2}\times I_{2}$, and $I_{3}\times I_{3}$ covariance matrices, respectively. The Kronecker structure is adopted to achieve dimensionality reduction. Without this structure, sampling $\boldsymbol{\Sigma}$ would require estimating $\mathcal{O}(I^{2})$ free parameters, making estimation unreliable with limited samples and computationally expensive—or even infeasible—as $I$ increases, especially in high-dimensional tensor time series. In contrast, imposing a Kronecker structure drastically reduces the number of free parameters to $\sum_{i=1}^{3}{I_{i}^{2}}$ and lowers the computational cost to $\mathcal{O}(\max(I_{1},I_{2},I_{3})^{6})$, thereby making the estimation more reliable and computationally tractable even in high-dimensional settings. To sample $\boldsymbol{\Sigma}_{i}$, we first compute the mode-$i$ matricization of the tensor $\mathcal{E}_{t}$, denoted by $\mathbf{E}_{t,(i)}$. It can be readily shown that $\mathbf{E}_{t,(i)}$ satisfies the following property:
@@ 5.3 Interpretation for Response Factor Time Series
In the BTAR-TK model with shrinkage priors, we adopt a dynamic factor interpretation for both responses and predictors. Specifically, the 6464 $(=R_{1}R_{2}R_{3}=R_{4}R_{5}R_{6})$ response and predictor factor time series, $\{\widetilde{\mathbf{B}}_{m}^{\prime}\mathbf{y}_{t}\}$ and $\{\widetilde{\mathbf{B}}^{\prime}\mathbf{y}_{t-1}\}$, are constructed; in the main text we focus on the response factors. Although the model is estimated on disaggregated country- and commodity-level data, most dimension-reduced response factors closely track more aggregated global indicators; the corresponding indicators are also displayed in the figure. Because each column of $\mathbf{B}_{i}$ is identified only up to sign, the sign of each factor is indeterminate; we orient each factor by choosing the sign that minimizes the RMSE with the corresponding index.
@@ 6 Conclusion
In summary, this paper develops a flexible and computationally efficient Bayesian framework for third-order tensor autoregressions using the Tucker decomposition. The main methodological contributions include incorporating stochastic volatility and hierarchical shrinkage priors into high-dimensional tensor-on-tensor regressions, and extending the commonly used CP decomposition to the more general Tucker structure. Extensive MCMC experiments demonstrate that when the true data-generating process is low-rank, the Tucker model consistently outperforms both CP-based approaches and standard Bayesian VARs. When the data-generating process is sparse, models equipped with multiway stick-breaking shrinkage priors exhibit superior performance relative to those without shrinkage. Moreover, even when the coefficient tensor does not favor the assumed low-rank structure, our approach remains competitive with Bayesian VAR benchmarks on small and medium datasets. Finally, in an empirical application to a $(19\times 19\times 15)$ dataset of multi-category bilateral trade flows, we uncover the principal countries and commodity clusters underlying global trade dynamics and provide a factor interpretation that links these latent factors to key global indicators.